OpenStudy (anonymous):

what is the limit of (3-x^2)/(x-3) as x approaches 3?

OpenStudy (helder_edwin):

\[ \large \lim_{x\to3}\frac{3-x^2}{x-3}=\frac{3-3^2}{3-3}=\frac{-6}{0}= \]

OpenStudy (anonymous):

the answer is dne? when i graphed it it was -1 the thing is is that i am not supposed to use a graphing calculator to solve it

OpenStudy (helder_edwin):

yes. it doesn't exist.